A teacher has a class consisting of 8 BOYS and 4 GIRLS. How many groups of 5 can he make if:
-in every group there should be AT LEAST 1 girl?..
Thanks so much in advance. I'm not really sure how to solve for this... ^_^ God Bless
please send me the answer ASAP! :D it's our assignment thank you!

Found 2 solutions by stanbon, sudhanshu_kmr:
Answer by stanbon(75887)

A teacher has a class consisting of 8 BOYS and 4 GIRLS. How many groups of 5 can he make if:
-in every group there should be AT LEAST 1 girl?..
-------
P(at least one) = 1 - P(no girls)
---
P(no girls) = 8C5/12C5 = 0.071
--
P(at least one girl) = 1-0.071 = 0.929..
============
Cheers,
Stan H.

Answer by sudhanshu_kmr(1152)


No. of ways to form group of 5 from total 12 students= 12C5 (no restriction)

no. of ways to form group of 5 from 8 boys DISABLED_event_only= 8C5 ........ ( no girl )

now, we subtract it from total possible no.

no. of ways to form group in which at least 1 girl = 12C5 - 8C5

= 792 - 56 = 736